OpenStudy (anonymous):

please help..integrate (16-19x^2)^3/2 /x^6 dx.. thank you.

OpenStudy (blockcolder):

Is that really 19x^2 or is it just 9x^2?

OpenStudy (anonymous):

sorry it is 9x^2

OpenStudy (blockcolder):

Let \(\large x=\frac{4}{3}\sin\theta\). Can you change all of the x's to theta?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

this topic was integration by parts

OpenStudy (blockcolder):

Can't do it by integration by parts. :(

OpenStudy (anonymous):

it is integration by substitution using trigonometric substitution..please help

OpenStudy (blockcolder):

\[\begin{align} \text{Let }&x=\frac{4}{3}\sin\theta\\ &dx=\frac{4}{3}\cos\theta d\theta\\ &x^6=\frac{2^{12}}{3^6}\sin^6\theta\\ &(16-9x^2)^{3/2}=8\cos^3\theta \end{align}\] Substitute the last three expressions into the integrand and simplify until you can evaluate the integral.

OpenStudy (anonymous):

thank u..i'm trying now..

OpenStudy (anonymous):

@blockcolder is there other way to solve this? because i was confuse where did u get the 4/3?thank u

OpenStudy (anonymous):

how and why does trig come into this?

OpenStudy (anonymous):